OpenStudy (anonymous):

solve (u+7)^3/2=-2 where u is a real number

OpenStudy (anonymous):

\[ (u+7)^{3/2}=-2 \rightarrow (u+7)=(-2)^{2/3}\] i.e. \[ (u+7)=\sqrt[3]{(-2)^2} \rightarrow (u+7)=\sqrt[3]{4} \rightarrow u=\sqrt[3]{4}-7\]

OpenStudy (anonymous):

If you want to find the value of u, u=-7+3√4....u=0